# Functors on the category of finite sets revisited

### v2

@@ 6.2. The image of the norm-like map
Using Corollary 6.7 it is straightforward to determine the kernel of this map.
@@
Considered as a $\mathfrak{S}_{n+1}$-module, $\Lambda^{n}(\overline{P})(\mathbf{n+1})$ identifies as the sign representation $\mathrm{sgn}_{n+1}$, by Lemma 5.7. To conclude, one uses that $\mathbbm{k}{\mathbf{FS}}(\mathbf{n+1},\mathbf{n})$, considered as a $\mathfrak{S}_{n+1}$-module, does not contain $\mathrm{sgn}_{n+1}$, which is shown by a straightforward verification. ∎
@@ 6.3. Evaluation of the simples
In the first case, $C_{\lambda}(\mathbf{t})\cong\mathbbm{k}{\mathbf{FI}}(\mathbf{n},\mathbf{t})\otimes_{\mathfrak{S}_{n}}S_{\lambda}$ as a $\mathbbm{k}\mathfrak{S}_{t}$-module. This is zero if $t<n$. If $t\geq n$, the permutation $\mathbbm{k}(\mathfrak{S}_{n}^{\mathrm{op}}\times\mathfrak{S}_{t})$-module $\mathbbm{k}{\mathbf{FI}}(\mathbf{n},\mathbf{t})$ is isomorphic to $\mathbbm{k}\mathfrak{S}_{t}/\mathfrak{S}_{t-n}\cong\mathbbm{k}\mathfrak{S}_{t}\otimes_{\mathfrak{S}_{t-n}}\mathsf{triv}_{t-n}$, using the residual right $\mathfrak{S}_{n}$-action (cf. Proposition 2.5). Thus one has the isomorphism of $\mathbbm{k}\mathfrak{S}_{t}$-modules
@@
$S_{\lambda}(P^{\mathbf{FA}})\cong\left\{\begin{array}[]{ll}\bigoplus_{\begin{subarray}{c}\nu\preceq\lambda\\ \lambda/\nu\in\mathrm{HS}\end{subarray}}S_{\nu}(\overline{P})&\lambda\neq(s,1^{n-s})\mbox{ for $1\leq s\leq n$}\\ \bigoplus_{\begin{subarray}{c}\nu\preceq\lambda\\ \nu\not\in\{(1^{n-s+1}),(1^{n-s})\}\\ \lambda/\nu\in\mathrm{HS}\end{subarray}}S_{\nu}(\overline{P})\ \oplus\ \Lambda^{n-s+1}(P^{\mathbf{FA}})&\lambda=(s,1^{n-s})\mbox{ for $1\leq s\leq n$}.\end{array}\right.$
@@
The right hand side is calculated by Proposition 6.14 (this is either 00 or $1$, depending on the value of $\nu$). Since $n$ is positive, if this is non-zero we must have $|\nu|>0$. This yields the decomposition in the statement, using the identification of the projective covers given in Theorem 6.6. ∎
@@ 7.1. Introducing the projective cover of $\overline{P}^{\otimes n}$
$0\rightarrow\overline{P}\rightarrow P^{\mathbf{FA}}\rightarrow\overline{\mathbbm{k}}=\overline{P}^{\otimes 0}\rightarrow 0$
@@
$0\rightarrow\Lambda^{n+1}(\overline{P})\rightarrow\mathbb{P}_{n}\stackrel{{\scriptstyle\mathfrak{p}_{n}}}{{\rightarrow}}\overline{P}^{\otimes n}\rightarrow 0$
@@
The first statement follows from the fact that $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{n+1}(\overline{P}),\overline{P}^{\otimes n})=0$, a consequence of Theorem 4.4. The second follows from the projectivity of $\mathbb{P}_{n}$ together with the fact that $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{n},\Lambda^{n+1}(\overline{P}))=0$, since $\Lambda^{n+1}(\overline{P})$ is not a composition factor of the cosocle of $\mathbb{P}_{n}$. This is seen as follows: by construction the cosocle of $\mathbb{P}_{n}$ identifies with that of $\overline{P}^{\otimes n}$; the latter does not contain $\Lambda^{n+1}(\overline{P})$ as a composition factor. ∎
@@ 7.2. Projective generators and the associated Morita equivalence
where, on the right, $S_{\lambda}$ is considered as a right $\mathfrak{S}_{n}^{\mathrm{op}}$-module. Such identifications are used henceforth without further comment.
@@
A standard Morita equivalence argument gives:
@@
In this part we implement the results of Part II. In particular, Theorem 9.5, which shows how to calculate the composition factors of a $\mathbbm{k}\mathbf{FA}$-module $F$, is applied notably to the cases $F=\overline{P}^{\otimes n}$ and its projective cover. These results are applied in Theorem 11.9 to calculate the space of morphisms between the $\mathbbm{k}\mathbf{FA}$-modules $\overline{P}^{\otimes m}$ and $\overline{P}^{\otimes n}$, for $m,n\in\mathbb{N}$. This extends the partial calculation Theorem 4.4.
@@
8. Inverting $-\odot\mathsf{triv}$
Working in the category of $\mathbbm{k}\mathbf{FB}$-modules, one has the functor $-\odot\mathsf{triv}$, where $\mathsf{triv}$ is as in Notation 2.4. The purpose of this section is to explain that this functor is invertible in a suitable sense (see Proposition 8.3 for the basic statement). This also applies when working with $\mathbbm{k}\mathbf{FB}^{\mathrm{op}}$-modules, mutatis mutandis.
@@ 8.1. The basic result
We require to work with isomorphism classes of objects and with virtual objects; to do so we pass to the Grothendieck group of $\mathbbm{k}\mathbf{FB}$-modules. This is possible, since all the $\mathbbm{k}\mathbf{FB}$-modules considered here have the property that, evaluated on any finite set, they have finite dimension. Such technical details are suppressed, for simplicity of exposition.
@@
    For $M,N$ $\mathbbm{k}\mathbf{FB}$-modules, denote by $[M]\odot[N]$ the class $[M\odot N]$. This extends bilinearly to define $\odot$ on the whole Grothendieck group.
For example, one has $[\mathsf{triv}]$ in the Grothendieck group and the associated operation $-\odot[\mathsf{triv}]$, corresponding to the functor $-\odot\mathsf{triv}$. For the sign representations, we work with the infinite sum
@@
This is a standard result; for completeness, a proof is given. When $n=0$, there is only one term that appears, namely the trivial representation. For $0<n\in\mathbb{N}$, consider the evaluation on $\mathbf{n}$, thus passing to the Grothendieck group of $\mathbbm{k}\mathfrak{S}_{n}$-modules. The term on the left hand side of (2.3) yields
@@
Using the identification given in Example 2.6, one obtains the equality $[\mathbbm{k}{\mathbf{FS}}]=\mathcal{S}(0)\odot_{\mathbf{FB}}[\mathbbm{k}\mathbf{FA}]$ in the Grothendieck ring of $\mathbbm{k}\mathbf{FB}$-bimodules.
@@ 8.2. A truncated variant
where $S_{(n-k,1^{k})}$ is understood to be 00 for $n=k$. Now, for each $n\geq k$, $S_{(n-k+1,1^{k-1})}\oplus S_{(n-k,1^{k})}\cong\mathsf{sgn}_{k}\odot\mathsf{triv}_{n-k}$, by the Pieri rule (as in the proof of Proposition 8.3). The sum $\bigoplus_{n\geq k}\mathsf{sgn}_{k}\odot\mathsf{triv}_{n-k}$ identifies with $\mathsf{sgn}_{k}\odot\mathsf{triv}$, by the definition of the latter. ∎
The invertibility of $-\odot\mathsf{triv}$ yields the following:
@@
$[H(k)]\odot\mathcal{S}(0)=\mathcal{S}(k).$
@@
$[H(k)]\odot\mathcal{S}(0)\ +\ [H(k+1)]\odot\mathcal{S}(0)=[\mathsf{sgn}_{k}\odot\mathsf{triv}]\odot\mathcal{S}(0)=[\mathsf{sgn}_{k}],$
using Proposition 8.3 for the second equality, since $\odot$ is associative. Now, by the inductive hypothesis, we have $[H(k)]\odot\mathcal{S}(0)=\mathcal{S}(k)$, hence
$[H(k+1)]\odot\mathcal{S}(0)=[\mathsf{sgn}_{k}]-\mathcal{S}(k).$
@@
The identity of Proposition 8.9 is equivalent to the equality $[H(k)]=\mathcal{S}(k)\odot[\mathsf{triv}]$ in the Grothendieck group of $\mathbbm{k}\mathbf{FB}$-modules. This can be viewed as a ‘truncated’ version of Proposition 8.3.
@@ 9. Calculating $\mathrm{Hom}(\mathbb{P}_{\bullet},F)$
$P^{\mathbf{FA}}_{\bullet}$ ::: $\mathbf{n}\mapsto P^{\mathbf{FA}}_{\mathbf{n}}$
$\overline{P}^{\otimes\bullet}$ ::: $\mathbf{n}\mapsto\overline{P}^{\otimes n}.$
@@ 9.1. Splitting off projective covers of exterior power functors
$\overline{P}^{\otimes\bullet}/\Lambda$ ::: $\mathbf{n}\mapsto\overline{P}^{\otimes n}/\Lambda^{n}(\overline{P}).$
@@
$0\rightarrow(P^{\mathbf{FA}}_{\bullet}/\Lambda)(\mathbf{n})\rightarrow P^{\mathbf{FA}}_{\mathbf{n}}\rightarrow P^{\mathbf{FA}}_{\Lambda,\mathbf{n}}\rightarrow 0$
@@ 9.2. How to calculate $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},-)$
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},F)]=[\overline{F}]\odot_{\mathbf{FB}}\mathcal{S}(0)\quad+\quad\sum_{k\geq 1}(\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}F(\mathbf{k}))\otimes\mathcal{S}(k-1),$
@@
As in Example 7.8, it is straightforward to calculate $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{n+1}(P^{\mathbf{FA}}),F)$. The key step is to calculate the $\mathbbm{k}\mathfrak{S}_{n}$-module $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n}/\Lambda^{n}(\overline{P}),F).$
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes\bullet}/\Lambda,F)]=[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\bullet}/\Lambda,F)]\odot_{\mathbf{FB}}\mathcal{S}(0).$
Now, since $[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\bullet},F)]=[F]$ by Yoneda’s lemma, this gives
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\bullet}/\Lambda,F)]=[F]-[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\Lambda,\bullet},F)].$
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\Lambda,\bullet},F)]=[F(0)\otimes\mathsf{triv}_{0}]+\sum_{n>0}\sum_{k=1}^{n}[\big{(}\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}F(\mathbf{k})\big{)}\otimes S_{(n-k+1,1^{k-1})}].$
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\Lambda,\bullet},F)]=\sum_{k\geq 0}[\big{(}\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}F(\mathbf{k})\big{)}\otimes H(k)],$
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes\bullet}/\Lambda,F)]=[F]\odot_{\mathbf{FB}}\mathcal{S}(0)\quad-\quad\sum_{k\geq 0}\big{(}\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}F(\mathbf{k})\big{)}\otimes\mathcal{S}(k).$
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes\bullet}/\Lambda,F)]=[\overline{F}]\odot_{\mathbf{FB}}\mathcal{S}(0)\quad-\quad\sum_{k\geq 1}\big{(}\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}F(\mathbf{k})\big{)}\otimes\mathcal{S}(k).$
It remains to calculate $[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},F)]$ by reintegrating the contribution from
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{\bullet+1}(P^{\mathbf{FA}}),F)$$\cong$ $\bigoplus_{m\geq 0}(\mathsf{sgn}_{m+1}\otimes_{\mathfrak{S}_{m+1}}F(\mathbf{m+1}))\ \otimes\ \mathsf{sgn}_{m}=\bigoplus_{k\geq 1}(\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}F(\mathbf{k}))\ \otimes\ \mathsf{sgn}_{k-1}.$

### v3

@@ 6.2. The image of the norm-like map
Using Corollary 6.7 it is straightforward to determine the kernel of this map, as follows.
@@
Considered as a $\mathfrak{S}_{n+1}$-module, $\Lambda^{n}(\overline{P})(\mathbf{n+1})$ identifies as the sign representation $\mathrm{sgn}_{n+1}$, by Lemma 5.6. To conclude, one uses that $\mathbbm{k}{\mathbf{FS}}(\mathbf{n+1},\mathbf{n})$, considered as a $\mathfrak{S}_{n+1}$-module, does not contain $\mathrm{sgn}_{n+1}$, which is straightforward. ∎
@@ 6.3. Evaluation of the simples
In the first case, $C_{\lambda}(\mathbf{t})\cong\mathbbm{k}{\mathbf{FI}}(\mathbf{n},\mathbf{t})\otimes_{\mathfrak{S}_{n}}S_{\lambda}$ as a $\mathbbm{k}\mathfrak{S}_{t}$-module. This is zero if $t<n$. If $t\geq n$, the permutation $\mathbbm{k}(\mathfrak{S}_{n}^{\mathrm{op}}\times\mathfrak{S}_{t})$-module $\mathbbm{k}{\mathbf{FI}}(\mathbf{n},\mathbf{t})$ is isomorphic to $\mathbbm{k}\mathfrak{S}_{t}/\mathfrak{S}_{t-n}\cong\mathbbm{k}\mathfrak{S}_{t}\otimes_{\mathfrak{S}_{t-n}}\mathsf{triv}_{t-n}$, using the residual right $\mathfrak{S}_{n}$-action (cf. the proof of Proposition 2.5). Thus one has the isomorphism of $\mathbbm{k}\mathfrak{S}_{t}$-modules
@@
$S_{\lambda}(P^{\mathbf{FA}})\cong\left\{\begin{array}[]{ll}\bigoplus_{\begin{subarray}{c}\nu\preceq\lambda\\ \lambda/\nu\in\mathrm{HS}\end{subarray}}S_{\nu}(\overline{P})&\lambda\neq(s,1^{n-s})\mbox{ for $1\leq s\leq n$}\\ \Lambda^{n-s+1}(P^{\mathbf{FA}})\oplus\bigoplus_{\begin{subarray}{c}\nu\preceq\lambda\\ \nu\not\in\{(1^{n-s+1}),(1^{n-s})\}\\ \lambda/\nu\in\mathrm{HS}\end{subarray}}S_{\nu}(\overline{P})&\lambda=(s,1^{n-s})\mbox{ for $1\leq s\leq n$}.\end{array}\right$.
@@
The right hand side is calculated by Proposition 6.14 (this is either 0 or $1$, depending on the value of $\nu$). Since $n$ is positive, if this is non-zero we must have $|\nu|>0$. This yields the decomposition in the statement, using the identification of the projective covers given in Theorem 6.6. ∎
@@ 7.1. Introducing the projective cover of $\overline{P}^{\otimes n}$
$0\rightarrow\overline{P}\rightarrow P^{\mathbf{FA}}\rightarrow\overline{\mathbbm{k}}=\overline{P}^{\otimes 0}\rightarrow 0.$
@@
$0\rightarrow\Lambda^{n+1}(\overline{P})\rightarrow\mathbb{P}_{n}\stackrel{{\scriptstyle\mathfrak{p}_{n}}}{{\rightarrow}}\overline{P}^{\otimes n}\rightarrow 0.$
@@
The first statement follows from the fact that $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{n+1}(\overline{P}),\overline{P}^{\otimes n})=0$, a consequence of Corollary 4.21. The second follows from the projectivity of $\mathbb{P}_{n}$ together with the fact that $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{n},\Lambda^{n+1}(\overline{P}))=0$, since $\Lambda^{n+1}(\overline{P})$ is not a composition factor of the head of $\mathbb{P}_{n}$ (the largest semi-simple quotient of $\mathbb{P}_{n}$, sometimes called the cosocle). This is seen as follows: by construction the head of $\mathbb{P}_{n}$ identifies with that of $\overline{P}^{\otimes n}$; the latter does not contain $\Lambda^{n+1}(\overline{P})$ as a composition factor. ∎
@@ 7.2. Projective generators and the associated Morita equivalence
where, on the right, $S_{\lambda}$ is considered as a right $\mathfrak{S}_{n}$-module. Such identifications are used henceforth without further comment.
@@
Since $\{\mathbb{P}_{n}\mid n\in\mathbb{N}\cup\{-1\}\}$ is a set of small projective generators of $\mathcal{F}(\mathbf{FA})$, Freyd’s theorem (see [3, Theorem 3.1] for a version over $\mathbb{Z}$) gives:
@@
In this part we implement the results of Part II. In particular, Theorem 9.5, which shows how to calculate the composition factors of a $\mathbbm{k}\mathbf{FA}$-module $F$, is applied notably to the cases $F=\overline{P}^{\otimes n}$ and its projective cover. These results are applied in Theorem 11.9 to calculate the space of morphisms between the $\mathbbm{k}\mathbf{FA}$-modules $\overline{P}^{\otimes m}$ and $\overline{P}^{\otimes n}$, for $m,n\in\mathbb{N}$. This extends the partial calculation Corollary 4.21.
@@
8. Inverting $\odot\mathsf{triv}$
Working in the category of $\mathbbm{k}\mathbf{FB}$-modules, one has the functor $\odot\mathsf{triv}$, where $\mathsf{triv}$ is as in Notation 2.4. The purpose of this section is to explain that this functor is invertible in a suitable sense (see Proposition 8.3 for the basic statement). This also applies when working with $\mathbbm{k}\mathbf{FB}^{\mathrm{op}}$-modules, mutatis mutandis.
@@ 8.1. The basic result
We require to work with isomorphism classes of objects and with virtual objects; to do so we pass to the Grothendieck group of $\mathbbm{k}\mathbf{FB}$-modules. This is possible, since all the $\mathbbm{k}\mathbf{FB}$-modules considered here have the property that, evaluated on a finite set, they have finite dimension. Such technical details are suppressed, for simplicity of exposition.
@@
    For $M,N$ $\mathbbm{k}\mathbf{FB}$-modules, denote by $M]\odot[N]$ the class $M\odot N]$. This extends bilinearly to define $\odot$ on the whole Grothendieck group.
For example, one has $\mathsf{triv}]$ in the Grothendieck group and the associated operation $\odot[\mathsf{triv}]$, corresponding to the functor $\odot\mathsf{triv}$. For the sign representations, we work with the infinite sum
@@
This is a standard result; for completeness, a proof is given. Consider the evaluation on $\mathbf{n}$, thus passing to the Grothendieck group of $\mathbbm{k}\mathfrak{S}_{n}$-modules. When $n=0$, there is only one term that appears, namely the trivial representation. For $0<n\in\mathbb{N}$, the term on the left hand side of (2.3) yields
@@
Using the identification given in Example 2.6, one obtains the equality $\mathbbm{k}{\mathbf{FS}}]=\mathcal{S}(0)\odot_{\mathbf{FB}}[\mathbbm{k}\mathbf{FA}]$ in the Grothendieck ring of $\mathbbm{k}\mathbf{FB}$-bimodules.
@@ 8.2. A truncated variant
where $S_{(n-k,1^{k})}$ is understood to be 0 for $n=k$. Now, for each $n\geq k$, $S_{(n-k+1,1^{k-1})}\oplus S_{(n-k,1^{k})}\cong\mathsf{sgn}_{k}\odot\mathsf{triv}_{n-k}$, by the Pieri rule (as in the proof of Proposition 8.3). The sum $\bigoplus_{n\geq k}\mathsf{sgn}_{k}\odot\mathsf{triv}_{n-k}$ identifies with $\mathsf{sgn}_{k}\odot\mathsf{triv}$, by the definition of the latter. ∎
The invertibility of $\odot\mathsf{triv}$ yields the following:
@@
$H(k)]\odot\mathcal{S}(0)=\mathcal{S}(k).$
@@
$H(k)]\odot\mathcal{S}(0)\ +\ [H(k+1)]\odot\mathcal{S}(0)=[\mathsf{sgn}_{k}\odot\mathsf{triv}]\odot\mathcal{S}(0)=[\mathsf{sgn}_{k}],$
using Proposition 8.3 for the second equality, since $\odot$ is associative. Now, by the inductive hypothesis, we have $H(k)]\odot\mathcal{S}(0)=\mathcal{S}(k)$, hence
$H(k+1)]\odot\mathcal{S}(0)=[\mathsf{sgn}_{k}]-\mathcal{S}(k).$
@@
The identity of Proposition 8.9 is equivalent to the equality $H(k)]=\mathcal{S}(k)\odot[\mathsf{triv}]$ in the Grothendieck group of $\mathbbm{k}\mathbf{FB}$-modules. This can be viewed as a ‘truncated’ version of Proposition 8.3.
@@ 9. Calculating $\mathrm{Hom}(\mathbb{P}_{\bullet},F)$
$P^{\mathbf{FA}}_{\bullet}$ $:$ $\mathbf{n}\mapsto P^{\mathbf{FA}}_{\mathbf{n}}$
$\overline{P}^{\otimes\bullet}$ $:$ $\mathbf{n}\mapsto\overline{P}^{\otimes n}.$
@@ 9.1. Splitting off projective covers of exterior power functors
$\overline{P}^{\otimes\bullet}/\Lambda$ $:$ $\mathbf{n}\mapsto\overline{P}^{\otimes n}/\Lambda^{n}(\overline{P}).$
@@
$0\rightarrow(P^{\mathbf{FA}}_{\bullet}/\Lambda)(\mathbf{n})\rightarrow P^{\mathbf{FA}}_{\mathbf{n}}\rightarrow P^{\mathbf{FA}}_{\Lambda,\mathbf{n}}\rightarrow 0.$
@@ 9.2. How to calculate $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},-)$
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},F)]=[\overline{F}]\odot_{\mathbf{FB}}\mathcal{S}(0)\quad+\quad\sum_{k\geq 1}(\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}F(\mathbf{k}))\otimes\mathcal{S}(k-1),$
@@
As in Example 7.8, it is straightforward to calculate $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{n+1}(P^{\mathbf{FA}}),F)$. The key step is thus to calculate the $\mathbbm{k}\mathfrak{S}_{n}$-module $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n}/\Lambda^{n}(\overline{P}),F).$
@@
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes\bullet}/\Lambda,F)]=[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\bullet}/\Lambda,F)]\odot_{\mathbf{FB}}\mathcal{S}(0).$
Now, since $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\bullet},F)]=[F]$ by Yoneda’s lemma, this gives
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\bullet}/\Lambda,F)]=[F]-[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\Lambda,\bullet},F)].$
@@
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\Lambda,\bullet},F)]=[F(0)\otimes\mathsf{triv}_{0}]+\sum_{n>0}\sum_{k=1}^{n}[\big{(}\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}F(\mathbf{k})\big{)}\otimes S_{(n-k+1,1^{k-1})}].$
@@
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\Lambda,\bullet},F)]=\sum_{k\geq 0}[\big{(}\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}F(\mathbf{k})\big{)}\otimes H(k)],$
@@
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes\bullet}/\Lambda,F)]=[F]\odot_{\mathbf{FB}}\mathcal{S}(0)\quad-\quad\sum_{k\geq 0}\big{(}\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}F(\mathbf{k})\big{)}\otimes\mathcal{S}(k).$
@@
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes\bullet}/\Lambda,F)]=[\overline{F}]\odot_{\mathbf{FB}}\mathcal{S}(0)\quad-\quad\sum_{k\geq 1}\big{(}\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}F(\mathbf{k})\big{)}\otimes\mathcal{S}(k).$
It remains to calculate $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},F)]$ by reintegrating the contribution from
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{\bullet+1}(P^{\mathbf{FA}}),F)$ $\cong$ $\bigoplus_{m\geq 0}(\mathsf{sgn}_{m+1}\otimes_{\mathfrak{S}_{m+1}}F(\mathbf{m+1}))\ \otimes\ \mathsf{sgn}_{m}$
$=$ $\bigoplus_{k\geq 1}(\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}F(\mathbf{k}))\ \otimes\ \mathsf{sgn}_{k-1}.$